# Liu and Zhang

### v2

@@
[…]
Liu and Zhang
[…]
PCL for RMAB with General Observation Models
[…]
PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models
[…]
@@
[…]
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic operations research problem that involves a learner making choices among actions with uncertain/random rewards in order to maximize the expected return in long-run. The MAB problem is often associated with the important exploration-exploitation tradeoff and was initially proposed by Robbins (1952). Starting from the classic stochastic scheduling problem, the ongoing development of various MAB models makes it applicable to a wide range of practical fields, including clinical trials, recommendation systems, cognitive communications, and financial investments (Gittins 1979, Berry and Fristedt 1985, Press 2009, Farias and Madan 2011, Hoffman et al. 2011, Shen et al. 2015).
@@
In the classic Bayesian MAB problem, the bandit machine has a total of $N$ arms and one player pulling an arm in each time slot. After a player selected one of these arms and activated it, a random reward will be accrued depending on the activated arm and its current state. In this process, all states are visible, and only the state of the arm being activated changes according to a Markov chain. The goal of the MAB problem is to search a policy that maximizes the long-term cumulative reward. Gittins (1979) proved that the classic MAB problem can be solved optimally by an index policy, referred to as Gittins index, and the player only needs to activate the arm with the largest index at each moment. The index policy significantly reduces the complexity of the problem from being exponential with the number of arms to being linear. After that, Whittle (1988) extended the MAB problem to the restless MAB (RMAB) model, in which $K$ arms can be activated at each moment, and the arms not chosen may also undergo state transitions over time. However, for the general RMAB problem with a finite state space, Papadimitriou and Tsitsiklis (1994) showed that the computational complexity of finding the optimal policy is already PSPACE-hard. Fortunately, Whittle (1988) generalized Gittins index to Whittle index that provides a possible solution to RMAB by considering Lagrangian relaxation and duality. Accordingly, Whittle index is optimal under a relaxed constraint that the time-average number of activated arms is $K$. For the original problem, the Whittle index policy has been proven to be asymptotically optimal per-arm wise as the number of arms goes to infinity under certain conditions (Weber and Weiss 1990, 1991). Nonetheless, not every RMAB is indexable, i.e., Whittle index may not exist. And the establishment of indexability is itself a difficult problem. Even if indexability is proven for a particular RMAB, analytical solutions of the Whittle index function can still be hard to obtain. For some important categories of RMAB models, Whittle indexability and its strong performance have been demonstrated in the literature, such as the dual-speed bandit problem and partially observable RMAB (Glazebrook et al. 2002, Ahmad et al. 2009, Liu and Zhao 2010, Gittins et al. 2011, Liu 2021, Liu et al. 2024).
@@
In this paper, we extend the partially observable RMAB considered in Liu et al. (2010, 2024) to general observation models. The previous work only considers special classes of observation errors and noises within the framework of partially observation Markov decision processes (POMDP) (Zhao and Sadler 2007, Liu et al. 2010, 2024). But general algorithms for POMDP often suffer from curse of dimensionality and becomes cumbersome when the value functions for dynamic programming are too complex (Smallwood and Sondik 1973, Sondik 1978). In order to deal with general observation models, we need an alternative methodology to design efficient algorithms for systems with large sets of parameters. In the interdisciplinary field of operations research and stochastic optimization, an analytical method called “Achievable Region” emerged in the 1990s. Interestingly, this method transforms a time-series optimization problem into linear programming (LP) problems that can be solved efficiently (Coffman and Mitrani 1980, Federgruen and Groenevelt 1988a, b, Gelenbe and Mitrani 2010). The challenge of using this method is mainly on proving the feasibility of such transformation (Bertsimas 1995). Referred to as the general conservation law (GCL), Bertsimas and Niño-Mora (1996) proposed the required structure of performance measures for the achievable region to be a polyhedron called extended polymatroid. Based on GCL, they extended Klimov’s algorithm to several stochastic scheduling problems, including the classic MAB problem (Klimov 1975, Thomas 1991). By relaxing the GCL restrictions, Niño-Mora (2001) proposed the partial conservation law (PCL) and specified conditions for an RMAB to satisfy PCL and subsequently lead to the numerical calculation of Whittle index. Later on, Niño-Mora (2002) offered an economic explanation of PCL-indexability and its relation to Whittle indexability where Whittle index can be interpreted as the optimal marginal cost rate. A more comprehensive summary of this work can be found in Niño-Mora (2007). However, all such work considers finite state spaces except that Frostig and Weiss (1999) extended the GCL framework for classic MAB to countable state spaces. In this paper, we extend the PCL framework to the RMAB with a general parametrization of observation model and an infinite state space. The main challenge of this extension is due to the high-dimensional probability state space (belief states) with complex dynamics for transitions over time. By establishing the weak duality between LP formulations, we build the PCL framework for analyzing our RMAB and subsequently design an efficient algorithm for calculating Whittle index when PCL-indexability is satisfied. Finally, we demonstrate the superior performance of our algorithm by numerical simulations.
@@ 2.1 Model Formulation
In this section, we begin with the formulation of the RMAB problem with general observation models. Assume the system has $N$ arms and the state space of the $n$th arm is $\mathcal{M}_{n}$. Therefore the entire system state space is $\mathcal{M}=\prod_{n=1}^{N}\mathcal{M}_{n}$. At each moment, the actual state of each arm undergos a state transition based on its own Markov probability transition matrix, and we will select $K$ arms to activate. For those activated arms, we can observe their states (with errors) and accrue some reward dependent on the observed states and the actual states. For those arms that are not activated, we can neither observe their states nor obtain reward from them, and their states still transit over time. Let $S_{n}(t),O_{n}(t)\in\mathbb{Z}^{+}$ be respectively the true and observed states of arm $n$ in slot $t$ and $A(t)$ the set of activated arms in slot $t$. First consider a single-arm process (i.e. $N=1$ and $\mathcal{M}=\mathcal{M}_{n}$) and drop the subscript $n$ (sometimes the time index $t$ is also dropped if referring to the same time index) for convenience. Suppose the state transition matrix, error matrix and reward matrix are $P=\{p_{ij}\}_{\mathcal{M}\times\mathcal{M}}$, $E=\{\varepsilon_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ and $R=\{r_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ respectively. For error matrix, $\varepsilon_{ij}$ represents the probability that the observed state is $j$ when the true state is $i$, that is, $\varepsilon_{ij}=P(O=j|S=i)$, while $r_{ij}$ represents the reward obtained when the true state is $i$ and the observed state is $j$ (different observations may lead to different sub-actions and thus different rewards following). Denoted by ${\Omega}_{a}=\{(\omega_{1},\cdots,\omega_{M})|\sum_{i=1}^{M}\omega_{i}=1,\omega_{i}\geq 0,1\leq i\leq M:=|\mathcal{M}|\}$ the belief space, for each $\boldsymbol{\omega}(t)=(\omega_{1}(t),\cdots,\omega_{M}(t))\in{\Omega_{a}}$, $\omega_{i}(t)$ represents the conditional probability that the true state is $i$ in slot $t$ (based on past observations). If the current belief state is $(\omega_{1},\cdots,\omega_{M})$, the current expected reward is $\sum_{i=1}^{M}\omega_{i}(\sum_{j=1}^{M}\varepsilon_{ij}r_{ij})$. In addition, we may sometimes receive additional feedback related to the true state and the immediate reward obtained that helps us better estimate the actual arm state (e.g., ACK/NAK in a communication channel to indicate whether the data was successfully transmitted (Liu et al. 2024)). Assume that there are $L$ feedback states (positive integers) in total that encompass all possible observations, and denote the feedback state at time $t$ by $F(t)~{}(L\leq M^{2})$. Define $\rho_{ij}=P(F(t)=j|S(t)=i)$. By Bayes rule, we have
@@ 2.2 Whittle Index
Definition 2.1 (Whittle Indexability)
@@
Definition 2.2 (Whittle Index)
@@ 2.3 Belief State Space
Definition 2.3 ($T$-step state space)
@@ 2.5 Partial Conservation Law
$T_{\boldsymbol{\omega}_{i}}^{\Omega}=\begin{cases}1+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}T_{\boldsymbol{\omega}_{i}}^{\Omega},&\boldsymbol{\omega}_{i}\in\Omega\\ \beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}T_{\boldsymbol{\omega}_{i}}^{\Omega}&\boldsymbol{\omega}_{i}\notin\Omega\end{cases}$ (28)
@@
$R_{\boldsymbol{\omega}_{i}}^{\Omega}=\begin{cases}R_{\boldsymbol{\omega}_{i}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}R_{\boldsymbol{\omega}_{i}}^{\Omega},&\boldsymbol{\omega}_{i}\in\Omega\\ \beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}R_{\boldsymbol{\omega}_{i}}^{\Omega}&\boldsymbol{\omega}_{i}\notin\Omega\end{cases}.$ (29)
@@
For the auxiliary arm, we have $A_{0}^{\{0\}}=1$ and $W_{0}^{\{0\}}=\lambda$. Generally, we can extend these variables to the multi-arm case. Assume that there are two arms and their state spaces are $\Omega(\boldsymbol{\omega})$ and $\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ with $\Omega(\boldsymbol{\omega})\cap\Omega^{\prime}(\boldsymbol{\omega}^{\prime})=\emptyset$. Then for $\Omega\subset\Omega(\boldsymbol{\omega})$, $\Omega^{\prime}\subset\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ and $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega})$, we define $T_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=T_{\boldsymbol{\omega_{i}}}^{\Omega},R_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=R_{\boldsymbol{\omega_{i}}}^{\Omega},A_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=A_{\boldsymbol{\omega_{i}}}^{\Omega}$ and $W_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=W_{\boldsymbol{\omega_{i}}}^{\Omega}$. Previously, Niño-Mora provided the relationships among $T_{\boldsymbol{\omega_{i}}}^{\Omega}$, $R_{\boldsymbol{\omega_{i}}}^{\Omega}$, $A_{\boldsymbol{\omega_{i}}}^{\Omega}$, and $W_{\boldsymbol{\omega_{i}}}^{\Omega}$ for finite-state problems in (Niño-Mora 2001, 2002). Here, we prove that these conservation relations also hold for countable state spaces:
Theorem 2.4 (Decomposition Law)
@@
Proof 2.5
Proof. Assume the initial belief state is $\boldsymbol{\omega_{k}}$. For convenience, abbreviate $x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{k}})$ and $x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{k}})$ as $x_{\boldsymbol{\omega_{i}}}^{1}$ and $x_{\boldsymbol{\omega_{i}}}^{0}$. In this problem, we also have the constraint equation:
@@
[…]
We can also perform certain deformations on two equations. Note that $x_{0}^{1}(\pi,\boldsymbol{\omega_{j}})=\frac{1}{1-\beta}-T_{\boldsymbol{\omega_{j}}}^{\pi}$, equation (36) can be written as
@@
Definition 2.6 (Partial Conservation Law (PCL))
@@
Definition 2.7 (Full subset chain of $\tilde{\Omega}$)
@@
Condition (i) in Definition 2.8 characterizes part of the boundary of the achievable region through the constraints between performance measures for an infinite dimensional linear programming problem, and condition (ii) means that the objective function can take its maximum value at a unique vertex in the achievable region while the policy corresponding to this vertex is a state priority policy obtained by sorting according to the size of the index $\lambda_{\boldsymbol{\omega}}$. Next, we show that if the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then we can construct a unique full subset chain of $\tilde{\Omega}_{0}:=\tilde{\Omega}\cup\{\{0\}\cup\Omega,\Omega\in\tilde{\Omega}\}$. Assume $\exists!C\in\mathcal{C}(\tilde{\Omega})$, s.t.$\forall\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$ with $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, we have $\lambda_{\boldsymbol{\omega_{i}}}^{C}<\lambda_{\boldsymbol{\omega_{j}}}^{C}$. We can construct the full subset chain $C^{\prime}$ of $\tilde{\Omega}_{0}$ as follows:
@@
Without loss of generality, here we can assume $\bigcup\limits_{\boldsymbol{\omega}:\lambda_{\boldsymbol{\omega}}\leq\lambda}\Omega_{\boldsymbol{\omega}}^{C}\in\tilde{\Omega}$, otherwise we just need to add it to $\tilde{\Omega}$. Under such a construction, it is easy to verify that $C^{\prime}$ is a full subset chain of $\tilde{\Omega}_{0}$, and it satisfies the second condition of Definition 2.8. Furthermore, we have the following characterization of the unique full subset chain of $\tilde{\Omega}_{0}$:
Lemma 2.9
@@
Proof 2.10
[…]
By combining Theorem 2.4 and Lemma 2.9, the following relationship can be established:
Lemma 2.11
@@
Proof 2.12
Proof. First we consider the case that $\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$. Let $\Omega=\Omega_{\boldsymbol{\omega_{i}}},\pi=\pi_{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}$ in Theorem 2.4, then
@@
[…]

### v3

# PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models

@@
Abstract
@@
1School of Mathematics and Physics, Xi’an Jiaotong-Liverpool University, Suzhou, 215123, Jiangsu, China.
2Jiangsu National Center for Applied Mathematics, Nanjing, 210093, Jiangsu, China.
3Bank of Communications, Shanghai, 200120, China.
Corresponding author. E-mail(s): Keqin.Liu@xjtlu.edu.cn;
Keywords Restless bandit  $\cdot$ POMDP  $\cdot$ Countable state space  $\cdot$ Partial conservation law  $\cdot$ Whittle index
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic operations research problem that involves a learner making choices among actions with uncertain/random rewards in order to maximize the expected return in long-run. The MAB problem is often associated with the important exploration-exploitation tradeoff and was initially proposed by Robbins (1952). Starting from the classic stochastic scheduling problem, the ongoing development of various MAB models makes it applicable to a wide range of practical fields, including clinical trials, recommendation systems, cognitive communications, and financial investments (Gittins, 1979; Berry and Fristedt, 1985; Press, 2009; Farias and Madan, 2011; Hoffman et al., 2011; Shen et al., 2015).
@@
In the classic Bayesian MAB problem, the bandit machine has a total of $N$ arms and one player pulling an arm in each time slot. After a player selected one of these arms and activated it, a random reward will be accrued depending on the activated arm and its current state. In this process, all states are visible, and only the state of the arm being activated changes according to a Markov chain. The goal of the MAB problem is to search a policy that maximizes the long-term cumulative reward. Gittins (1979) proved that the classic MAB problem can be solved optimally by an index policy, referred to as Gittins index, and the player only needs to activate the arm with the largest index at each moment. The index policy significantly reduces the complexity of the problem from being exponential with the number of arms to being linear. After that, Whittle (1988) extended the MAB problem to the restless MAB (RMAB) model, in which $K$ arms can be activated at each moment, and the arms not chosen may also undergo state transitions over time. However, for the general RMAB problem with a finite state space, Papadimitriou and Tsitsiklis (1994) showed that the computational complexity of finding the optimal policy is already PSPACE-hard. Fortunately, Whittle (1988) generalized Gittins index to Whittle index that provides a possible solution to RMAB by considering Lagrangian relaxation and duality. Accordingly, Whittle index is optimal under a relaxed constraint that the time-average number of activated arms is $K$. For the original problem, the Whittle index policy has been proven to be asymptotically optimal per-arm wise as the number of arms goes to infinity under certain conditions (Weber and Weiss, 1990a, b). Nonetheless, not every RMAB is indexable, i.e., Whittle index may not exist. And the establishment of indexability is itself a difficult problem. Even if indexability is proven for a particular RMAB, analytical solutions of the Whittle index function can still be hard to obtain. For some important categories of RMAB models, Whittle indexability and its strong performance have been demonstrated in the literature, such as the dual-speed bandit problem and partially observable RMAB (Glazebrook et al., 2002; Ahmad et al., 2009; Liu and Zhao, 2010; Gittins and Weber, 2011; Liu, 2025; Liu et al., 2024).
@@
In this paper, we extend the partially observable RMAB considered in Liu et al. (2010, 2024) to general observation models. The previous work only considers special classes of observation errors and noises within the framework of partially observation Markov decision processes (POMDP) (Zhao and Sadler, 2007; Liu et al., 2010, 2024). But general algorithms for POMDP often suffer from curse of dimensionality and becomes cumbersome when the value functions for dynamic programming are too complex (Smallwood and Sondik, 1973; Sondik, 1978). In order to deal with general observation models, we need an alternative methodology to design efficient algorithms for systems with large sets of parameters. In the interdisciplinary field of operations research and stochastic optimization, an analytical method called “Achievable Region” emerged in the 1990s. Interestingly, this method transforms a time-series optimization problem into linear programming (LP) problems that can be solved efficiently (Coffman Jr and Mitrani, 1980; Federgruen and Groenevelt, 1988a; Federgruen and Groenevelt, 1988b; Gelenbe and Mitrani, 2010). The challenge of using this method is mainly on proving the feasibility of such transformation (Bertsimas, 1995). Referred to as the general conservation law (GCL), Bertsimas and Niño-Mora (1996) proposed the required structure of performance measures for the achievable region to be a polyhedron called extended polymatroid. Based on GCL, they extended Klimov’s algorithm to several stochastic scheduling problems, including the classic MAB problem (Klimov, 1975; Division and Tsoucas, 1991). By relaxing the GCL restrictions, Niño-Mora (2001) proposed the partial conservation law (PCL) and specified conditions for an RMAB to satisfy PCL and subsequently lead to the numerical calculation of Whittle index. Later on, Niño-Mora (2002) offered an economic explanation of PCL-indexability and its relation to Whittle indexability where Whittle index can be interpreted as the optimal marginal cost rate. A more comprehensive summary of this work can be found in Niño-Mora (2007). However, all such work considers finite state spaces except that Frostig and Weiss (1999) extended the GCL framework for classic MAB to countable state spaces. In this paper, we extend the PCL framework to the RMAB with a general parametrization of observation model and an infinite state space. The main challenge of this extension is due to the high-dimensional probability state space (belief states) with complex dynamics for transitions over time. By establishing the weak duality between LP formulations, we build the PCL framework for analyzing our RMAB and subsequently design an efficient algorithm for calculating Whittle index when PCL-indexability is satisfied. Finally, we demonstrate the superior performance of our algorithm by numerical simulations.
@@ 2.1 Model Formulation
In this section, we begin with the formulation of the RMAB problem with general observation models. Assume the system has $N$ arms and the state space of the $n$th arm is $\mathcal{M}_{n}$. Therefore the entire system state space is $\mathcal{M}=\prod_{n=1}^{N}\mathcal{M}_{n}$. At each moment, the actual state of each arm undergos a state transition based on its own Markov probability transition matrix, and we will select $K$ arms to activate. For those activated arms, we can observe their states (with errors) and accrue some reward dependent on the observed states and the actual states. For those arms that are not activated, we can neither observe their states nor obtain reward from them, and their states still transit over time. Let $S_{n}(t),O_{n}(t)\in\mathbb{Z}^{+}$ be respectively the true and observed states of arm $n$ in slot $t$ and $A(t)$ the set of activated arms in slot $t$. First consider a single-arm process (i.e. $N=1$ and $\mathcal{M}=\mathcal{M}_{n}$) and drop the subscript $n$ (sometimes the time index $t$ is also dropped if referring to the same time index) for convenience. Suppose the state transition matrix, error matrix and reward matrix are $P=\{p_{ij}\}_{\mathcal{M}\times\mathcal{M}}$, $E=\{\varepsilon_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ and $R=\{r_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ respectively. For error matrix, $\varepsilon_{ij}$ represents the probability that the observed state is $j$ when the true state is $i$, that is, $\varepsilon_{ij}=P(O=j|S=i)$, while $r_{ij}$ represents the reward obtained when the true state is $i$ and the observed state is $j$ (different observations may lead to different sub-actions and thus different rewards following). Denoted by ${\Omega}_{a}=\{(\omega_{1},\cdots,\omega_{M})|\sum_{i=1}^{M}\omega_{i}=1,\omega_{i}\geq 0,1\leq i\leq M:=|\mathcal{M}|\}$ the belief space, for each $\boldsymbol{\omega}(t)=(\omega_{1}(t),\cdots,\omega_{M}(t))\in{\Omega_{a}}$, $\omega_{i}(t)$ represents the conditional probability that the true state is $i$ in slot $t$ (based on past observations). If the current belief state is $(\omega_{1},\cdots,\omega_{M})$, the current expected reward is $\sum_{i=1}^{M}\omega_{i}(\sum_{j=1}^{M}\varepsilon_{ij}r_{ij})$. In addition, we may sometimes receive additional feedback related to the true state and the immediate reward obtained that helps us better estimate the actual arm state (e.g., ACK/NAK in a communication channel to indicate whether the data was successfully transmitted (Liu et al., 2024)). Assume that there are $L$ feedback states (positive integers) in total that encompass all possible observations, and denote the feedback state at time $t$ by $F(t)~{}(L\leq M^{2})$. Define $\rho_{ij}=P(F(t)=j|S(t)=i)$. By Bayes rule, we have
@@ 2.2 Whittle Index
Definition 1 (Whittle Indexability).
@@
Definition 2 (Whittle Index).
@@ 2.3 Belief State Space
Definition 3 ($T$-step state space).
@@ 2.5 Partial Conservation Law
$T_{\boldsymbol{\omega}_{i}}^{\Omega}=\begin{cases}1+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}T_{\boldsymbol{\omega}_{j}}^{\Omega},&\boldsymbol{\omega}_{i}\in\Omega\\ \beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}T_{\boldsymbol{\omega}_{j}}^{\Omega}&\boldsymbol{\omega}_{i}\notin\Omega\end{cases}$ (28)
@@
$R_{\boldsymbol{\omega}_{i}}^{\Omega}=\begin{cases}R_{\boldsymbol{\omega}_{i}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}R_{\boldsymbol{\omega}_{j}}^{\Omega},&\boldsymbol{\omega}_{i}\in\Omega\\ \beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}R_{\boldsymbol{\omega}_{j}}^{\Omega}&\boldsymbol{\omega}_{i}\notin\Omega\end{cases}.$ (29)
@@
For the auxiliary arm, we have $A_{0}^{\{0\}}=1$ and $W_{0}^{\{0\}}=\lambda$. Generally, we can extend these variables to the multi-arm case. Assume that there are two arms and their state spaces are $\Omega(\boldsymbol{\omega})$ and $\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ with $\Omega(\boldsymbol{\omega})\cap\Omega^{\prime}(\boldsymbol{\omega}^{\prime})=\emptyset$. Then for $\Omega\subset\Omega(\boldsymbol{\omega})$, $\Omega^{\prime}\subset\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ and $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega})$, we define $T_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=T_{\boldsymbol{\omega_{i}}}^{\Omega},R_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=R_{\boldsymbol{\omega_{i}}}^{\Omega},A_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=A_{\boldsymbol{\omega_{i}}}^{\Omega}$ and $W_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=W_{\boldsymbol{\omega_{i}}}^{\Omega}$. Previously, Niño-Mora provided the relationships among $T_{\boldsymbol{\omega_{i}}}^{\Omega}$, $R_{\boldsymbol{\omega_{i}}}^{\Omega}$, $A_{\boldsymbol{\omega_{i}}}^{\Omega}$, and $W_{\boldsymbol{\omega_{i}}}^{\Omega}$ for finite-state problems in (Niño-Mora, 2001, 2002). Here, we prove that these conservation relations also hold for countable state spaces:
Theorem 1 (Decomposition Law).
@@
Proof.
Assume the initial belief state is $\boldsymbol{\omega_{k}}$. For convenience, abbreviate $x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{k}})$ and $x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{k}})$ as $x_{\boldsymbol{\omega_{i}}}^{1}$ and $x_{\boldsymbol{\omega_{i}}}^{0}$. In this problem, we also have the constraint equation:
@@
$\Box$
We can also perform certain transformations on two equations. Note that $x_{0}^{1}(\pi,\boldsymbol{\omega_{j}})=\frac{1}{1-\beta}-T_{\boldsymbol{\omega_{j}}}^{\pi}$, equation (36) can be written as
@@
Definition 4 (Partial Conservation Law (PCL)).
@@
Definition 5 (Full subset chain of $\tilde{\Omega}$).
@@
Condition (i) in Definition 6 characterizes part of the boundary of the achievable region through the constraints between performance measures for an infinite dimensional linear programming problem, and condition (ii) means that the objective function can take its maximum value at a unique vertex in the achievable region while the policy corresponding to this vertex is a state priority policy obtained by sorting according to the size of the index $\lambda_{\boldsymbol{\omega}}$. Next, we show that if the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then we can construct a unique full subset chain of $\tilde{\Omega}_{0}:=\tilde{\Omega}\cup\{\{0\}\cup\Omega,\Omega\in\tilde{\Omega}\}$. Assume $\exists!C\in\mathcal{C}(\tilde{\Omega})$, s.t.$\forall\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$ with $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, we have $\lambda_{\boldsymbol{\omega_{i}}}^{C}<\lambda_{\boldsymbol{\omega_{j}}}^{C}$. We can construct the full subset chain $C^{\prime}$ of $\tilde{\Omega}_{0}$ as follows:
@@
Without loss of generality, here we can assume $\bigcup\limits_{\boldsymbol{\omega}:\lambda_{\boldsymbol{\omega}}\leq\lambda}\Omega_{\boldsymbol{\omega}}^{C}\in\tilde{\Omega}$, otherwise we just need to add it to $\tilde{\Omega}$. Under such a construction, it is easy to verify that $C^{\prime}$ is a full subset chain of $\tilde{\Omega}_{0}$, and it satisfies the second condition of Definition 6. Furthermore, we have the following characterization of the unique full subset chain of $\tilde{\Omega}_{0}$:
Lemma 2.
@@
Proof.
If $\Omega_{\boldsymbol{\omega_{i}}}$ has the form of (45), then it clearly satisfies the definition of the full subset chain. We will prove that the full subset chain that satisfies condition (ii) in Definition 6 is exactly in the form of (45). Assume that there exists $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ such that $\lambda_{\boldsymbol{\omega_{j}}}>\lambda_{\boldsymbol{\omega_{i}}}$, so that $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{i}}}$. Since $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{j}}}$, by condition (ii) in Definition 5, $\Omega_{\boldsymbol{\omega_{j}}}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}$. By condition (ii) in Definition 6, we have $\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}$, this contradicts with $\lambda_{\boldsymbol{\omega_{j}}}>\lambda_{\boldsymbol{\omega_{i}}}$. Similarly, assume that there exists $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ such that $\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}$, but $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}$, then $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{j}}}$ implies that $\Omega_{\boldsymbol{\omega_{i}}}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}$, which contradicts with the hypothesis. $\Box$
By combining Theorem 1 and Lemma 2, the following relationship can be established:
Lemma 3.
@@
Proof.
First we consider the case that $\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$. Let $\Omega=\Omega_{\boldsymbol{\omega_{i}}},\pi=\pi_{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}$ in Theorem 1, then
@@
If $\boldsymbol{\omega_{j}}=0$, then we have $A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}=A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash{\{\boldsymbol{\omega_{i}}\}}}=1$ and $W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash{\{\boldsymbol{\omega_{i}}\}}}=\lambda$; if $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ and $\boldsymbol{\omega_{i}}=0$, then by definition, $A_{\boldsymbol{\omega_{j}}}^{\Omega_{0}}=A_{\boldsymbol{\omega_{j}}}^{\Omega_{0}\backslash{\{0\}}}$ and $W_{\boldsymbol{\omega_{j}}}^{\Omega_{0}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{0}\backslash{\{0\}}}$. In both cases, the lemma is clearly true. $\Box$